Determine the answer to the following equation with the correct number of significant figures: (17.000 + 2.30) × 1.0100 =
1
Expert's answer
2020-09-16T05:11:31-0400

QUESTION #133397

Determine the answer to the following equation with the correct number of significant figures: (17.000 + 2.30) × 1.0100 =

ANSWER

(17.000 + 2.30) × 1.0100

=19.30×1.0100

=19.493

=19.5 (3s.f)